# Distributed Discrete-time Dynamic Outer Approximation of the Intersection of Ellipsoids

### v1

@@ V-B Constant inputs
for all $i\in\mathcal{I}$. Moreover, there exists $f^{*}_{1},\dots,f^{*}_{\textsf{N}}\in\mathbb{R}$ such that
$\lim_{k\to\infty}f(\mathbf{Q}_{i}[k])=f^{*}_{i}$
@@
Let $\lambda_{j}^{i}=1$ if $j=i$ and $\lambda_{j}^{i}=0$ otherwise, and $\lambda_{\mathbf{P}}^{i}=0$. Therefore,
$\lambda_{\mathbf{P}}^{i}\mathbf{P}_{i}[k]^{-1}+\sum_{j\in\mathcal{N}_{i}}\lambda_{j}^{i}\mathbf{Q}_{j}[k-1]=\mathbf{Q}_{i}[k-1]$
and $\lambda_{\mathbf{P}}^{i}+\sum_{j\in\mathcal{N}_{i}}\lambda_{j}^{i}=1.$ Hence, $\mathbf{Q}_{i}[k-1]\in\mathcal{C}_{i}[k]$. Therefore, (3) follows by noting that both $\mathbf{Q}_{i}[k],\mathbf{Q}_{i}[k-1]\in\mathcal{C}_{i}[k]$ but that $\mathbf{Q}_{i}[k]$ is the minimizer of $f(\bullet)$ over $\mathcal{C}_{i}[k]$.
For the last part of the lemma, note that $\mathcal{C}^{*}$ is closed and convex due to Lemma 2-1) such that item 4) of Assumption 2 implies $f$ is lower bounded over $\mathcal{C}^{*}$. Moreover, since $\mathcal{C}_{i}\subseteq\mathcal{C}^{*}$ then $f$ attains the same lower bound over $\mathcal{C}_{i}$. Furthermore, combine (3) with such lower bound to conclude that $f^{*}_{i}:=\lim_{k\to\infty}f(\mathbf{Q}_{i}[k])$ must exist due to the monotone convergence theorem [39, Theorem 3.14]. ∎
Lemma 4.
Let Assumptions 1 and 2 hold and ${\bar{\theta}}=1$. Moreover, let $\mathcal{G}$ be connected. Then, there exists $\mathbf{Q}_{{\textsf{eq}}}\in\mathbb{S}_{+}^{n}$ such that $\lim_{k\to\infty}\mathbf{Q}_{i}[k]=\mathbf{Q}_{{\textsf{eq}}}$ for all $i\in\mathcal{I}$.
@@
Proof.
We study the equilibrium of (2) as follows. Let
$\ell=\operatorname*{arg\,min}_{i\in\mathcal{I}}f_{i}^{*}$
with $f_{i}^{*}$ taken from Lemma 3. Moreover, let
$\mathcal{N}_{\ell}^{r}=\{i\in\mathcal{I}:i\in\mathcal{N}_{j},j\in\mathcal{N}_{\ell}^{r-1}\}$
and $\mathcal{N}_{\ell}^{0}=\mathcal{N}_{\ell}$. This is, $\mathcal{N}_{\ell}^{r-1}$ contains the set of neighbors of neighbors of $\ell\in\mathcal{I}$ after $r$ hops. We now show that $f_{\ell}^{*}=f_{i}^{*}$ for all $i\in\mathcal{N}_{\ell}^{r}$ and all $r\in\{0,1,\dots\}$ by induction, for some $k$ in which equilibrium is attained. For the induction base, note that there exists $\mathbf{Q}^{*}_{\ell}\in\mathbb{S}_{+}^{n}$ such that $f(\mathbf{Q}^{*}_{\ell})=f_{\ell}^{*}\leq f_{i}^{*}\leq f(\mathbf{Q})$ for all $\mathbf{Q}\in\mathcal{C}_{i}[k]$ with $i\in\mathcal{N}_{\ell}$. Thus, (2) implies $f_{i}^{*}=f(\mathbf{Q}_{\ell}^{*})$ since $\mathbf{Q}_{\ell}^{*}\in\mathcal{C}_{i}[k]$. The induction step follows in the same way. Finally, since $\mathcal{G}$ is connected, then $\mathcal{N}_{\ell}^{d}=\mathcal{I}$ where $d$ is the diameter of $\mathcal{G}$. Therefore, $f_{i}^{*}=f_{\ell}^{*}$ for all $i\in\mathcal{I}$. Finally, note that (2) being a convex program implies that $f(\mathbf{Q})=f_{\ell}^{*}$ is attained for a single $\mathbf{Q}\in\mathbb{S}_{+}^{n}$ which we denote as $\mathbf{Q}_{\textsf{eq}}$. ∎
Convergence to an equilibrium does not immediately imply convergence to the global optimum, since the feasible sets at each node do not coincide in general. Therefore, we need to prove that $\mathbf{Q}_{\mathsf{eq}}=\mathbf{Q}^{*}$.
@@
First, item 3) of Lemma 2 implies that in equilibrium, $\mathcal{C}_{i}\subseteq\mathcal{C}^{*}$, which are constant $\mathcal{C}_{i}=\mathcal{C}_{i}[k],\mathcal{C}^{*}=\mathcal{C}^{*}[k]$ due to Lemma 4. Now, assume $\mathbf{Q}^{*}\notin\mathcal{C}_{i}$ for all $i$ for a contradiction. Item 2) of Lemma 1 implies $\mathbf{Q}_{{\textsf{eq}}}\in{\textsf{rebdr}}(\mathcal{C}_{i})$. In equilibrium, it follows that
@@
1) ${\textsf{dim}}(\mathcal{P})=0$, with $\mathcal{P}$ consisting of a single point, namely $\mathbf{Q}^{*}=\mathbf{P}_{1}^{-1}=\dots=\mathbf{P}_{\textsf{N}}^{-1}$, which is a contradiction since from (2) it follows that $\mathbf{Q}_{i}[0]=\mathbf{Q}_{\textsf{eq}}=\mathbf{P}_{i}^{-1}$.
2) ${\textsf{dim}}(\mathcal{P})=1$ and $\mathbf{Q}_{\textsf{eq}}\in\mathcal{P}$. In this case, $\mathcal{P}$ necessarily consists of a line segment from $\mathbf{P}_{i}^{-1}$ and $\mathbf{P}_{j}^{-1}$ with $i\neq j$. Note that in this case, the line segment must comply $\mathcal{P}=\mathcal{S}$ for the star $\mathcal{S}$. However, this imply $f(\mathbf{Q}_{\textsf{eq}})\leq f(\mathbf{Q}^{*})$ which contradicts $\mathbf{Q}_{\textsf{eq}}\neq\mathbf{Q}^{*}$.
3) ${\textsf{dim}}(\mathcal{P})\geq 1$ and $\mathbf{Q}_{\textsf{eq}}\notin\mathcal{P}$. In this case, denote with $\mathcal{H}={\textsf{co}}(\mathcal{P}\cup\{\mathbf{Q}_{\textsf{eq}}\})$. Note that $\mathcal{H}$ is a polytope over the (possibly) lower dimensional space ${\textsf{span}}\{\mathbf{P}_{1}^{-1},\dots,\mathbf{P}_{\textsf{N}}^{-1},\mathbf{Q}_{\textsf{eq}}\}\subset\mathbb{S}_{+}^{n}$, with $\mathbf{Q}_{\textsf{eq}}$ as one corner, with at least two faces touching it due to the dimension of $\mathcal{P}$. Let $\mathcal{W}(\alpha):=\{\mathbf{Q}\in\mathbb{S}_{+}^{n}:f(\mathbf{Q})\leq\alpha\}$ the level sets of $f$, which are convex for any $\alpha\in\mathbb{R}$. Moreover, note that $\mathcal{W}(f(\mathbf{Q}_{\textsf{eq}}))\subseteq\mathcal{C}^{*}\cap\mathcal{H}$. Otherwise, there would exist smaller $\alpha<f(\mathbf{Q}_{\textsf{eq}})$ such that $\mathcal{W}(f(\mathbf{Q}_{\textsf{eq}}))\cap\mathcal{S}\neq\varnothing$, which contradicts the fact that $f(\mathbf{Q}_{\textsf{eq}})\leq f(\mathbf{Q})$ for all $\mathbf{Q}\in\mathcal{S}$. Using the previous fact in combination with $\mathbf{Q}_{\textsf{eq}}\in\mathcal{W}(f(\mathbf{Q}_{\textsf{eq}}))$ hence, $\mathcal{W}(f(\mathbf{Q}_{\textsf{eq}}))$ has at least two supporting hyper-planes at $\mathbf{Q}_{\textsf{eq}}$, namely the faces of $\mathcal{H}$ meeting at that point. However, $f$ is smooth by Assumption 2, which imply that $\mathcal{W}(f(\mathbf{Q}_{\textsf{eq}}))$ must have a single supporting hyperplane at all points (see [40, Theorems 23.3 and 25.1 ]), leading to a contradiction.
@@
Henceforth, a contradiction is reached in any case implying that $\mathbf{Q}^{*}\in\mathcal{C}_{i}$ for some $i\in\mathcal{I}$. Then, item 1) of Lemma 2 allows the usage of item 1) of Lemma 1, which implies $\mathbf{Q}_{i}^{*}=\operatorname*{arg\,min}_{\mathbf{Q}\in\mathcal{C}_{i}}f(\mathbf{Q})=\mathbf{Q}^{*}$, being $\mathbf{Q}_{i}^{*}=\mathbf{Q}_{{\textsf{eq}}}$ the unique equilibrium of (2) due to Lemma 4, completing the proof. ∎
@@ V-C Dynamic inputs
In this section, we provide a proof for Theorem 1 in the general case with dynamic inputs. We prove each item of the theorem in the following:
@@
$\mathbf{Q}_{i}[k]\kern-2.84544pt=\kern-2.84544pt\lambda_{\mathbf{P}}^{i}\mathbf{P}_{i}[k]^{-1}\kern-2.84544pt+\kern-2.84544pt\frac{1}{\overline{\theta}}\kern-2.84544pt\sum_{j\in\mathcal{N}_{i}}\kern-2.84544pt\lambda_{j}^{i}\mathbf{Q}_{j}[k]\kern 0.0pt\preceq\kern 0.0pt\underline{p}\bigg{(}\kern-2.84544pt\lambda_{\mathbf{P}}^{i}\kern-2.84544pt+\kern-2.84544pt\frac{1}{\overline{\theta}}\kern-2.84544pt\sum_{j\in\mathcal{N}_{i}}\kern-2.84544pt\lambda_{j}^{i}\bigg{)}\mathbf{I}\kern 0.0pt\preceq\kern 0.0pt\underline{p}\mathbf{I}$
Item 3): First, let $\varepsilon,\underline{\theta},\overline{\theta}$ be such that $[\underline{\theta},\overline{\theta}]\subseteq[1-\varepsilon,1+\varepsilon]$. Note that if $\varepsilon=0$, then $\overline{\theta}=\underline{\theta}=1$ which corresponds to the static case. We now show the existence of $K\geq 0,\varepsilon>0$ such that $\|\mathbf{Q}_{i}[K]-\mathbf{Q}^{*}[K]\|<\delta$ independently of the initial conditions, by repeated use of a continuity argument. We denote the trajectories of a nominal version of the algorithm (2) with constant inputs $\mathbf{P}_{i}[k]=\mathbf{P}_{i}[0]$ as $\mathbf{Q}_{i}^{\varepsilon=0}[k]$. Due to stability and optimally of the nominal system established in Lemma 5, for any $\delta_{1}>0$ there exist $K^{\prime}\geq 0$ such that $\|\mathbf{Q}_{i}^{\varepsilon=0}[K^{\prime}]-\mathbf{Q}^{*,\varepsilon=0}[K^{\prime}]\|<\delta_{1}$. Moreover, note that $(1-\varepsilon)^{K^{\prime}}\mathbf{P}_{i}[0]\preceq\mathbf{P}_{i}[K^{\prime}]\preceq(1+\varepsilon)^{K^{\prime}}\mathbf{P}_{i}[0]$ by Assumption 1. Henceforth, for any $\delta_{2}>0$, there exists $\varepsilon>0$ sufficiently small to make $\mathbf{P}_{i}[K^{\prime}]$ as close as desired to $\mathbf{P}_{i}[0]$ to make $d_{H}(\mathcal{C}^{*}[K^{\prime}],\mathcal{C}^{*,\varepsilon=0}[K^{\prime}])<\delta_{2}$, where $d_{H}$ denotes the Hausdorff distance and $\mathcal{C}^{*,\varepsilon=0}[K]$ the feasible set in (1) in the nominal case. Similarly, for any $\delta_{3}>0$ there exists $\varepsilon>0$ such that $d_{H}(\mathcal{C}_{i}[K^{\prime}],\mathcal{C}_{i}^{\varepsilon=0}[K^{\prime}])<\delta_{3}$. Henceforth, by picking $\delta_{1},\delta_{2},\delta_{3}$ appropriately, $\varepsilon>0$ exist to make $\mathbf{Q}_{i}[K^{\prime}]$ as close as desired to $\mathbf{Q}_{i}^{\varepsilon=0}[K^{\prime}]$ (with error related to $\delta_{3}$), to $\mathbf{Q}^{*,\varepsilon=0}[K^{\prime}]$ (with error related to $\delta_{1}$) and to $\mathbf{Q}^{*}[K^{\prime}]$ (with error related to $\delta_{2}$) in that order, allowing $\|\mathbf{Q}_{i}[K^{\prime}]-\mathbf{Q}^{*}[K^{\prime}]\|<\delta$ for some $\delta$.
@@
Assume that $\|\mathbf{Q}_{i}[k]-\mathbf{Q}^{*}[k]\|<\alpha$ holds for some arbitrary $k$ and $\alpha<\delta$. Now, we show $\|\mathbf{Q}_{i}[k+1]-\mathbf{Q}^{*}[k+1]\|<\alpha$. The property follows by making $\varepsilon$ and $\alpha>0$ sufficiently small by taking $\mathbf{Q}^{*}[k]$ as close as desired to $\mathbf{Q}^{*}[k+1]$ and $\mathbf{Q}_{i}[k]$ close to $\mathbf{Q}_{i}[k+1]$, using a similar reasoning as before. Hence, there exists $\tilde{\varepsilon}=\inf\varepsilon(\mathbf{Q}_{1}[k],\dots,\mathbf{Q}_{\textsf{N}}[k])>0$ and $\tilde{\alpha}=\inf\alpha(\mathbf{Q}_{1}[k],\dots,\mathbf{Q}_{\textsf{N}}[k])$ if the infimum is taken over the set of possible $\mathbf{Q}_{i}[k]$, which is compact by the boundedness property established in item 2) of Theorem 1.

### v2

@@ V-B Constant inputs
for all $i\in\mathcal{I},{k=1,2,\dots}$. Moreover, there exists ${f_{1},\dots,f_{\textsf{N}}}\in\mathbb{R}$ such that $\lim_{k\to\infty}f(\mathbf{Q}_{i}[k])={f_{i}}$.
@@
Let $\lambda_{j}^{i}=1$ if $j=i$ and $\lambda_{j}^{i}=0$ otherwise, and $\lambda_{\mathbf{P}}^{i}=0$. Therefore, $\lambda_{\mathbf{P}}^{i}\mathbf{P}_{i}[k]^{-1}+\sum_{j\in\mathcal{N}_{i}}\lambda_{j}^{i}\mathbf{Q}_{j}[k-1]=\mathbf{Q}_{i}[k-1]$ and $\lambda_{\mathbf{P}}^{i}+\sum_{j\in\mathcal{N}_{i}}\lambda_{j}^{i}=1.$ Hence, $\mathbf{Q}_{i}[k-1]\in\mathcal{C}_{i}[k]$. Therefore, (3) follows by noting that both $\mathbf{Q}_{i}[k],\mathbf{Q}_{i}[k-1]\in\mathcal{C}_{i}[k]$ but that $\mathbf{Q}_{i}[k]$ is the minimizer of $f(\bullet)$ over $\mathcal{C}_{i}[k]$.
For the last part of the lemma, note that $\mathcal{C}^{*}$ is closed and convex due to Lemma 2-1) such that item 4) of Assumption 2 implies $f$ is lower bounded over $\mathcal{C}^{*}$. Moreover, since $\mathcal{C}_{i}{[k]}\subseteq\mathcal{C}^{*}$ then $f$ attains the same lower bound over $\mathcal{C}_{i}{[k]}$. Furthermore, combine (3) with such lower bound to conclude that ${f_{i}}:=\lim_{k\to\infty}f(\mathbf{Q}_{i}[k])$ must exist due to monotonicity in (3). ∎
@@
First, we study the equilibrium of (2) as follows. Let $\ell=\operatorname*{arg\,min}_{i\in\mathcal{I}}f_{i}^{*}$ with $f_{i}^{*}$ taken from Lemma 3. Moreover, let $\mathcal{N}_{\ell}^{r}=\{i\in\mathcal{I}:i\in\mathcal{N}_{j},j\in\mathcal{N}_{\ell}^{r-1}\}$ and $\mathcal{N}_{\ell}^{0}=\mathcal{N}_{\ell}$. This is, $\mathcal{N}_{\ell}^{r-1}$ contains the set of neighbors of neighbors of $\ell\in\mathcal{I}$ after $r$ hops. We now show that ${f_{\ell}=f_{i}}$ for all $i\in\mathcal{N}_{\ell}^{r}$ and all $r\in\{0,1,\dots\}$ by induction, for some $k$ in which equilibrium is attained. For the induction base, note that there exists ${\mathbf{Q}_{\ell}^{\prime}}\in\mathbb{S}_{+}^{n}$ such that $f({\mathbf{Q}_{\ell}^{\prime}})={f_{\ell}}\leq{f_{i}}\leq f(\mathbf{Q})$ for all $\mathbf{Q}\in\mathcal{C}_{i}[k]$ with $i\in\mathcal{N}_{\ell}$. Thus, (2) implies ${f_{i}}=f({\mathbf{Q}_{\ell}^{\prime}})$ since ${\mathbf{Q}_{\ell}^{\prime}}\in\mathcal{C}_{i}[k]$. The induction step follows in the same way. Finally, since $\mathcal{G}$ is connected, then $\mathcal{N}_{\ell}^{d}=\mathcal{I}$ where $d$ is the diameter of $\mathcal{G}$. Therefore, ${f_{i}}={f_{\ell}}$ for all $i\in\mathcal{I}$. Finally, note that (2) being a strictly convex program implies that $f(\mathbf{Q})={f_{\ell}}$ is attained for a single $\mathbf{Q}\in\mathbb{S}_{+}^{n}$ which we denote as $\mathbf{Q}_{\textsf{eq}}$.
Moreover, item 3) of Lemma 2 implies that in equilibrium, $\mathcal{C}_{i}\subseteq\mathcal{C}^{*}$, which are constant $\mathcal{C}_{i}=\mathcal{C}_{i}[k],\mathcal{C}^{*}=\mathcal{C}^{*}[k]$ for $k\geq 1$. Now, assume $\mathbf{Q}^{*}\notin\mathcal{C}_{i}$ for all $i$ for a contradiction. Item 2) of Lemma 1 implies $\mathbf{Q}_{{\textsf{eq}}}\in{\textsf{rebdr}}(\mathcal{C}_{i})$. In equilibrium, it follows that
@@
1) ${\textsf{dim}}(\mathcal{P})=0$, with $\mathcal{P}$ consisting of a single point $\mathbf{Q}^{*}=\mathbf{P}_{1}^{-1}=\dots=\mathbf{P}_{\textsf{N}}^{-1}$. However, note that from (2) it follows that $\mathbf{Q}_{i}[k]=\mathbf{Q}_{\textsf{eq}}=\mathbf{P}_{i}^{-1}{=\mathbf{Q}^{*}}$ and thus $\mathbf{Q}^{*}\in\mathcal{C}_{i}$, which is a contradiction.
2) ${\textsf{dim}}(\mathcal{P})=1$ and $\mathbf{Q}_{\textsf{eq}}\in\mathcal{P}$. In this case, $\mathcal{P}$ necessarily consists of a line segment from $\mathbf{P}_{i}^{-1}$ and $\mathbf{P}_{j}^{-1}$ with $i\neq j$. Note that in this case, the line segment must comply $\mathcal{P}=\mathcal{S}$ for the star $\mathcal{S}$. However, this imply $f(\mathbf{Q}_{\textsf{eq}})\leq f(\mathbf{Q}^{*})$. Due to strong convexity of $f$, either $f(\mathbf{Q}_{\textsf{eq}})<f(\mathbf{Q}^{*})$ which is impossible, or $\mathbf{Q}^{*}=\mathbf{Q}_{\textsf{eq}}=\mathbf{Q}_{i}[k]$ which leads to a contradiction similar to case 1).
3) ${\textsf{dim}}(\mathcal{P})\geq 1$ and $\mathbf{Q}_{\textsf{eq}}\notin\mathcal{P}$. In this case, denote with $\mathcal{H}={\textsf{co}}(\mathcal{P}\cup\{\mathbf{Q}_{\textsf{eq}}\})$. Note that $\mathcal{H}$ is a polytope over the (possibly) lower dimensional space ${\textsf{span}}\{\mathbf{P}_{1}^{-1},\dots,\mathbf{P}_{\textsf{N}}^{-1},\mathbf{Q}_{\textsf{eq}}\}\subset\mathbb{S}_{+}^{n}$, with $\mathbf{Q}_{\textsf{eq}}$ as one corner, with at least two faces touching it due to the dimension of $\mathcal{P}$. Let $\mathcal{W}(\alpha):=\{\mathbf{Q}\in\mathbb{S}_{+}^{n}:f(\mathbf{Q})\leq\alpha\}$ the level sets of $f$, which are convex for any $\alpha\in\mathbb{R}$. Moreover, note that $\mathcal{W}(f(\mathbf{Q}_{\textsf{eq}}))\subseteq\mathcal{C}^{*}\cap\mathcal{H}$. Otherwise, there would exist smaller $\alpha<f(\mathbf{Q}_{\textsf{eq}})$ such that $\mathcal{W}(f(\mathbf{Q}_{\textsf{eq}}))\cap\mathcal{S}\neq\varnothing$, which contradicts the fact that $f(\mathbf{Q}_{\textsf{eq}})\leq f(\mathbf{Q})$ for all $\mathbf{Q}\in\mathcal{S}$. Using the previous fact in combination with $\mathbf{Q}_{\textsf{eq}}\in\mathcal{W}(f(\mathbf{Q}_{\textsf{eq}}))$ hence, $\mathcal{W}(f(\mathbf{Q}_{\textsf{eq}}))$ has at least two supporting hyper-planes at $\mathbf{Q}_{\textsf{eq}}$, namely the faces of $\mathcal{H}$ meeting at that point. However, $f$ is smooth by Assumption 2, which imply that $\mathcal{W}(f(\mathbf{Q}_{\textsf{eq}}))$ must have a single supporting hyperplane at all points (see [44, Theorems 23.3 and 25.1]), contradicting the previous fact that $\mathcal{W}(f(\mathbf{Q}_{\textsf{eq}}))$ has two supporting hyper-planes.
@@
Henceforth, a contradiction is reached in any case implying that $\mathbf{Q}^{*}\in\mathcal{C}_{i}$ for some $i\in\mathcal{I}$. Then, item 1) of Lemma 2 allows the usage of item 1) of Lemma 1, which implies $\mathbf{Q}_{i}^{*}=\operatorname*{arg\,min}_{\mathbf{Q}\in\mathcal{C}_{i}}f(\mathbf{Q})=\mathbf{Q}^{*}$, being $\mathbf{Q}_{i}^{*}=\mathbf{Q}_{{\textsf{eq}}}$ the unique equilibrium of (2). ∎
@@ V-C Dynamic inputs
In this section, we provide a proof for Theorem 1 in the general case with dynamic inputs.
@@
[…]
Item 3): First, let $\varepsilon,\underline{\theta},\overline{\theta}$ be such that $[\underline{\theta},\overline{\theta}]\subseteq[1-\varepsilon,1+\varepsilon]$. Note that if $\varepsilon=0$, then $\overline{\theta}=\underline{\theta}=1$ which corresponds to the static case. Hence, we expect that small $\delta>0$ will lead to a small $\varepsilon>0$, preventing it to be arbitrarily big. To quantify this, we now show the existence of $K\geq{1},\varepsilon>0$ such that $\|\mathbf{Q}_{i}[K]-\mathbf{Q}^{*}[K]\|<\delta$ independently of the initial conditions, by repeated use of a continuity argument. We denote the trajectories of a nominal version of the algorithm (2) with constant inputs $\mathbf{P}_{i}[k]=\mathbf{P}_{i}[0]$ as $\mathbf{Q}_{i}^{\varepsilon=0}[k],{k\geq 1}$. Due to stability and optimally of the nominal system established in Lemma 4, for any $\delta_{1}>0$ there exist $K^{\prime}\geq 0$ such that $\|\mathbf{Q}_{i}^{\varepsilon=0}[K^{\prime}]-\mathbf{Q}^{*,\varepsilon=0}[K^{\prime}]\|<\delta_{1}$. Moreover, note that $(1-\varepsilon)^{K^{\prime}}\mathbf{P}_{i}[0]\preceq\mathbf{P}_{i}[K^{\prime}]\preceq(1+\varepsilon)^{K^{\prime}}\mathbf{P}_{i}[0]$ by Assumption 1. Henceforth, for any $\delta_{2}>0$, there exists $\varepsilon>0$ sufficiently small to make $\mathbf{P}_{i}[K^{\prime}]$ as close as desired to $\mathbf{P}_{i}[0]$ to make $d_{H}(\mathcal{C}^{*}[K^{\prime}],\mathcal{C}^{*,\varepsilon=0}[K^{\prime}])<\delta_{2}$, where $d_{H}$ denotes the Hausdorff distance and $\mathcal{C}^{*,\varepsilon=0}[K]$ the feasible set in (1) in the nominal case. Similarly, for any $\delta_{3}>0$ there exists $\varepsilon>0$ such that $d_{H}(\mathcal{C}_{i}[K^{\prime}],\mathcal{C}_{i}^{\varepsilon=0}[K^{\prime}])<\delta_{3}$. Henceforth, by picking $\delta_{1},\delta_{2},\delta_{3}$ appropriately, $\varepsilon>0$ exist to make $\mathbf{Q}_{i}[K^{\prime}]$ as close as desired to $\mathbf{Q}_{i}^{\varepsilon=0}[K^{\prime}]$ (with error related to $\delta_{3}$), to $\mathbf{Q}^{*,\varepsilon=0}[K^{\prime}]$ (with error related to $\delta_{1}$) and to $\mathbf{Q}^{*}[K^{\prime}]$ (with error related to $\delta_{2}$) in that order, allowing $\|\mathbf{Q}_{i}[K^{\prime}]-\mathbf{Q}^{*}[K^{\prime}]\|<\delta$ for some $\delta$.
@@
Assume that $\|\mathbf{Q}_{i}[k]-\mathbf{Q}^{*}[k]\|<\alpha$ holds for some arbitrary ${k\geq 0}$ and $\alpha<\delta$. Now, we show $\|\mathbf{Q}_{i}[k+1]-\mathbf{Q}^{*}[k+1]\|<\alpha$. The property follows by making $\varepsilon$ and $\alpha>0$ sufficiently small by taking $\mathbf{Q}^{*}[k]$ as close as desired to $\mathbf{Q}^{*}[k+1]$ and $\mathbf{Q}_{i}[k]$ close to $\mathbf{Q}_{i}[k+1]$, using a similar reasoning as before. Hence, there exists $\tilde{\varepsilon}=\inf\varepsilon(\mathbf{Q}_{1}[k],\dots,\mathbf{Q}_{\textsf{N}}[k])>0$ and $\tilde{\alpha}=\inf\alpha(\mathbf{Q}_{1}[k],\dots,\mathbf{Q}_{\textsf{N}}[k])$ if the infimum is taken over the set of possible $\mathbf{Q}_{i}[k]$, which is compact by the boundedness property established in item 2) of Theorem 1.